# Non-linear degenerate parabolic flow equations and a finer differential structure on Wasserstein spaces

### v1

@@ 3.7.1 The linear equation and LAFAS
First, we introduce a linear analogon to (3.14).
Definition 3.39 (MDE).
Let $T>0$, $p\geq 2$ and $\mu_{0}\in Wc_{p}(M)$. Let $V:[0,T]\to\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$ be continuous with respect to the pp–Wasserstein distance. Let $0,T]\to\mathcal{P}(M)$, $t\mapsto\mu(t)$ be a path of measures. We say that $\mu(t)$ satisfies the linear measure differential equation (MDE) for $V$ with initial value $\mu_{0}$ if $\mu(0)=\mu_{0}$, $\mu$ is narrowly continuous and for all $s\in[0,T],$ and all $\varphi$ smooth and compactly supported on $M$, we have
$\int_{M}\varphi\operatorname{d}\mu(s)-\int_{M}\varphi\operatorname{d}\mu(0)=\int_{0}^{s}\int_{M}\square^{V[t]}\varphi\operatorname{d}\mu(t)\operatorname{d}t,$ (3.125)
@@
$\square^{v_{t}}\varphi=\int_{\mathcal{X}(M)}\frac{1}{2}\mathcal{L}_{X-\overline{V}[\mu(t)]}^{2}(\varphi)+\mathcal{L}_{X}(\varphi)\operatorname{d}V_{t}.$ (3.126)
We introduce an appropriate analogon of the average flow approximation series for the MDE (3.125).
@@
Let $T>0$, $\nu:[0,T]\to\mathcal{W}_{p}(M)$ be continuous with respect to the pp–Wasserstein distance, $p\geq 2$, $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $\delta>0$ and $\mathcal{}\mathfrak{P}(\delta)$ be a partition of $[0,T]$ with maximal step–length $\mathfrak{p}(\mathcal{}\mathfrak{P}(\delta))\leq\delta$. Let $N\coloneqq|\mathcal{}\mathfrak{P}(\delta)|$ and index the elements of $\mathcal{}\mathfrak{P}(\delta)$ in increasing order. Define
@@
for all $t\in[\frac{x_{l+1}+x_{l}}{2},x_{l}]$ and
$\mu^{\mathcal{}\mathfrak{P}(\delta)}(t)=\bigl{(}e^{V[\nu_{x_{l}}]}_{2(t-\frac{x_{l+1}+x_{l}}{2})}\circ f^{V[\nu_{x_{l}}]}_{x_{l+1}-x_{l}}\bigr{)}\bigl{(}\mu^{\mathcal{}\mathfrak{P}(\delta)}\bigl{(}\frac{x_{l+1}+x_{l}}{2}\bigr{)}\bigr{)},$
for all $t\in[x_{l+1},\frac{x_{l+1}+x_{l}}{2}]$.
@@
Let $p\geq 2$, $T>0$ and $V:[0,T]\to\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$ be continuous. A narrowly continuous curve $\nu:[0,T]\to\mathcal{W}_{p}(M)$ is called a smooth solution of (3.125) if it solves (3.125) for $V$ and there exists a sequence $(\mathcal{}\mathfrak{P}_{n})_{n\in\mathbb{N}}$ of partitions of $[0,T]$ with maximal step–length $\mathfrak{p}(\mathcal{}\mathfrak{P}_{n})\to 0$ such that for all $n\in\mathbb{N}$, the LAFAS $\mu^{\mathcal{}\mathfrak{P}_{n}}$ converges pointwise narrowly to $\nu$.
Now, we readily obtain an existence result for equation (3.125) which is analogous to the Generalized Central Limit Theorem 3.35.
@@
Let $T>0$ and $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $\bigl{(}\mathcal{}\mathfrak{P}(\frac{1}{N})\bigr{)}_{N\in\mathbb{N}}$ be a family of partitions of $[0,T]$. Let $\nu:[0,T]\to\mathcal{W}_{p}(M)$ be continuous with respect to the pp–Wasserstein distance and $V$ be a continuous probability vector field on $\mathcal{W}_{p}(M)$ for $p>2$. Assume that $V$ has uniformly bounded pp–th moments, e.g, there exists a constant $B>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $\mathcal{M}_{p}(V[\mu])<B$.
@@
    Assume that $(M,g)$ is of bounded geometry. Assume furthermore, that there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]$ is supported in the centered ball with radius $R$. Then, there exists a subsequence $\biggl{(}\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N_{k}})}\biggr{)}_{k\in\mathbb{N}}$ of the Linear Average Flow Approximation Series $\biggl{(}\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\biggr{)}_{N\in\mathbb{N}}$ with initial value $\mu_{0}$ (See Definition 3.40) which uniformly converges with respect to the pp–Wasserstein distance to a $\frac{1}{p}$–Hölder–continuous solution $\mu:[0,T]\to\mathcal{W}_{p}(M)$ of (3.125) for $t\mapsto V[\nu(t)]$ with initial value $\mu_{0}$.
@@
    Now, we only assume that there exists $q<p$ such that $V$ is continuous for the $\mathcal{W}_{q}$–distance. Then, there exists a subsequence $\biggl{(}\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N_{k}})}\biggr{)}_{k\in\mathbb{N}}$ of the Linear Average Flow Approximation Series with initial value $\mu_{0}$ which converges to a solution $\mu:[0,T]\to\mathcal{W}_{p}(M)$ of (3.125) $t\mapsto V[\nu(t)]$ with initial value $\mu_{0}$.
@@
Introducing the linear equation (3.125) allows us to derive the uniqueness of solutions of (3.14) which are limit of AFAS from the uniqueness of solutions of (3.125) which are limit of LAFAS using a fixed point argument. This argument is detailed in our second main result, Theorem 3.43.
Theorem 3.43.
Let $p\geq 2$, $T>0$ and $\mu_{0}\in\mathcal{W}_{p}(M)$. Assume:
  • $(M,g)$ is of bounded geometry.
  • $V$ is a vector–field probability which satisfies:
    • $V$ is Lipschitz with respect to the pp–Wasserstein distance.
    • There exist $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]$ is supported in the centered ball with radius $R$.
  • For all $\frac{1}{p}$–Hölder–continuous $\mu:[0,T]\to\mathcal{W}_{p}(M)$ with $\mu(0)=\mu_{0}$, the linear MDE (3.125) has a unique smooth solution (See Definition 3.41) for $\mu$ with initial value $\mu_{0}$ in $\mathcal{W}_{p}(M)$.
Then:
[…]
  • For any sequence of partitions of $[0,T]$, $\biggl{(}\mathcal{}\mathfrak{P}(\frac{1}{N})\biggr{)}_{N\in\mathbb{N}}$ with $\lim_{N\to\infty}\mathfrak{p}(\mathcal{}\mathfrak{P}(\frac{1}{N}))=0$, the sequence of Average Flow Approximation Series $\biggl{(}\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\biggr{)}_{N\in\mathbb{N}}$ with initial value $\mu_{0}$ converges uniformly to this unique smooth solution in $\mathcal{W}_{p}(M)$.
Proof.
Let $T>0$ and $\biggl{(}\mathcal{}\mathfrak{P}(\frac{1}{N})\biggr{)}_{N\in\mathbb{N}}$ be a sequence of partitions of $[0,T]$. Let $\mathcal{S}:\mathcal{C}^{\frac{1}{p}}([0,T],\mathcal{W}_{p}(M))\to\mathcal{C}^{\frac{1}{p}}([0,T],\mathcal{W}_{p}(M))$ be the map sending a $\frac{1}{p}$–Hölder–continuous path $\mu:[0,T]\to\mathcal{W}_{p}(M)$ to the unique smooth solution of the corresponding linear MDE (3.125) with initial value $\mu_{0}$. We will show that $\mathcal{S}$ is a contraction with respect to the $\mathcal{W}_{p}$–distance for $T$ small enough.
Let us first work with $(M,g)=(\mathbb{R}^{n},g^{eucl})$. Let $t\in[0,T]$ and $\varepsilon>0$. Let $c_{1},c_{2}\in\mathcal{C}^{\frac{1}{p}}([0,T],\mathcal{W}_{p}(M))$. Denote for all $N\in\mathbb{N}$, $(x_{N,i})_{i\in\llbracket 0,I_{N}-1\rrbracket}\coloneqq\mathcal{}\mathfrak{P}(\frac{1}{N})$ in increasing order, and the respective LAFAS with steps in $\mathcal{}\mathfrak{P}(\frac{1}{N})$ by $\mu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}$ and $\nu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}$. Note that, for all $\mu,\nu\in\mathcal{W}_{p}(M)$, $\mathcal{W}_{2}(V[\mu],V[\nu])\leq 2R$. Hence, (3.60) in Lemma 3.31 yields that there exists $N\in\mathbb{N}$—depending on $\mathcal{W}_{2}(c_{1},c_{2})_{\infty}$—, such that for all $n\geq\mathbb{N}$,
[…]
Now, let $i\in I_{n}$ and suppose that there exists $C>0$ such that,
$\mathcal{W}_{2}\bigl{(}\mu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i}),\nu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i})\bigr{)}^{2}\leq C^{2}\mathcal{W}_{2}(c_{1},c_{2})_{\infty}^{2}.$
Then,
$\mathcal{W}_{2}(f^{V[c_{1}(x_{n,i})]}_{x_{n,i+1}-x_{n,i}}(\mu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i})),f^{V[c_{2}(x_{n,i})]}_{x_{n,i+1}-x_{n,i}}(\nu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i})))^{2}$ (3.128)
$\qquad\leq\biggl{(}\mathcal{W}_{2}\bigl{(}\mu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i}),\nu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i})\bigr{)}^{2}$
$\qquad+2(x_{n,i+1}-x_{n,i})\biggl{(}9L^{2}\mathcal{W}_{2}(c_{1},c_{2})^{2}_{\infty}+3LRC\mathcal{W}_{2}(c_{1},c_{2})^{2}_{\infty}\biggr{)}$
$\qquad+4R^{2}\biggl{(}RC\mathcal{W}_{2}(c_{1},c_{2})_{\infty}+3L\mathcal{W}_{2}(c_{1},c_{2})_{\infty}\biggr{)}\frac{(x_{n,i+1}-x_{n,i})^{\frac{3}{2}}}{6}$
$\qquad+\frac{2R^{4}}{3}(x_{n,i+1}-x_{n,i})^{2}\biggr{)}e^{8R^{2}(x_{n,i+1}-x_{n,i})}.$
In particular, for all $C^{2}\leq\mathcal{B}$ for some $\mathcal{B}>0$, there exists $M\geq N$—depending on $\mathcal{W}_{2}(c_{1},c_{2})_{\infty}$—such that for all $n\geq M$,
$\mathcal{W}_{2}(f^{V[c_{1}(x_{n,i})]}_{x_{n,i+1}-x_{n,i}}(\mu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i})),f^{V[c_{2}(x_{n,i})]}_{x_{n,i+1}-x_{n,i}}(\nu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i})))^{2}$ (3.129)
$\qquad\leq\biggl{(}C^{2}+2(x_{n,i+1}-x_{n,i})\bigl{(}9L^{2}+3LRC+8R^{2}C^{2}+\varepsilon\bigr{)}\biggr{)}\mathcal{W}_{2}(c_{1},c_{2})^{2}_{\infty}$
$\qquad\leq\biggl{(}C^{2}+2(x_{n,i+1}-x_{n,i})\bigl{(}9L^{2}+3LR\sqrt{\mathcal{B}}+8R^{2}\mathcal{B}+\varepsilon\bigr{)}\biggr{)}\mathcal{W}_{2}(c_{1},c_{2})^{2}_{\infty}.$
Now, denote by $t\mapsto\mu(t)$ and $t\mapsto\nu(t)$ the smooth solutions of (3.125) for $c_{1}$ and $c_{2}$ respectively. By assumption, the linear MDEs (3.125) for $c_{1}$ and $c_{2}$ have unique smooth solutions. Thus, for any sequence of partitions $\biggl{(}\mathcal{}\mathfrak{P}(\frac{1}{N})\biggr{)}_{N\in\mathbb{N}}$, the Linear Average Flow Approximation Series $\biggl{(}\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\biggr{)}_{N\in\mathbb{N}}$ and $\biggl{(}\nu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\biggr{)}_{N\in\mathbb{N}}$ converge uniformly to those solutions with respect to the pp–Wasserstein distance by Theorem 3.42. It follows by induction from (3.129) that there exist $T,\mathcal{K}>0$, such that for all $t\in[0,T]$,
$\mathcal{W}_{2}(\mu(t),\nu(t))\leq\mathcal{K}\sqrt{t}\mathcal{W}_{2}(c_{1},c_{2})_{\infty}.$ (3.130)
In particular,
$\mathcal{W}_{2}(\mu(t),\nu(t))_{\infty}\leq\mathcal{K}\sqrt{T}\mathcal{W}_{2}(c_{1},c_{2})_{\infty}.$ (3.131)
Therefore, choosing $T$ small enough, ensures that $\mathcal{S}$ is a contraction. thus, Banach’s fixed point theorem yields that $\mathcal{S}$ has a unique fixed point $s:[0,T]\to\mathcal{W}_{p}(M)$. Hence, (3.14) has a unique solution $s:[0,T]\to\mathcal{W}_{2}(M)$ with initial value $\mu_{0}$ on $[0,T]$, which is a limit of Linear Average Flow Approximation Series associated to $s$.
Let us conclude by showing that $s$ is also the unique solution of 3.14 in $\mathcal{W}_{p}(M)$ with initial value $\mu_{0}$, which is a limit of Average Flow Approximation Series. Suppose there exist two such solutions $s_{1}$ and $s_{2}$. Hence, there exist $\biggl{(}\mathcal{}\mathfrak{P}_{1}(\frac{1}{N})\biggr{)}_{N\in\mathbb{N}}$ and $\biggl{(}\mathcal{}\mathfrak{P}_{2}(\frac{1}{N})\biggr{)}_{N\in\mathbb{N}}$, two sequences of partitions of $[0,T]$ and Average Flow Approximation Series $s^{\mathcal{}\mathfrak{P}_{1}(\frac{1}{N})}_{1}$ and $s^{\mathcal{}\mathfrak{P}_{2}(\frac{1}{N})}_{2}$ converging uniformly in $\mathcal{W}_{p}(M)$ respectively to $s_{1}$ and $s_{2}$ with respective steps given by $\mathcal{}\mathfrak{P}_{1}(\frac{1}{N})$ and $\mathcal{}\mathfrak{P}_{2}(\frac{1}{N})$. Let $\varepsilon>0$. Since the convergence is uniform and $V$ is LL–Lipschitz with respect to the pp–Wasserstein distance, there exists $N_{0}\in\mathbb{N}$, such that for all $n\geq N_{0}$,
Now, it suffices to show that this implies that the LAFAS $\mathfrak{s}^{\mathcal{}\mathfrak{P}_{1}(\frac{1}{N})}_{1}$ and $\mathfrak{s}^{\mathcal{}\mathfrak{P}_{2}(\frac{1}{N})}_{2}$ associated to the linear MDEs (3.125) for $s_{1}$ and $s_{2}$ also converge to $s_{1}$ and $s_{2}$ respectively. This follows again by induction using (3.132) and (3.56) in Lemma 3.31. Thus, $s_{1}$ and $s_{2}$ are both fixed points of $\mathcal{S}$ and, hence, $s_{1}=s_{2}$.
@@ 3.7.2 Grönwall type estimates and Hausdorff–continuity of the set of solutions
Now, we prove two Grönwall type estimates for AFAS and LAFAS respectively for different initial values and identical initial values, but different vector–field probabilities.
@@
  • $V$ is a vector–field probability which satisfies:
    • There exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]$ is supported in the centered ball with radius $R$.
@@
Let $\varepsilon>0$. Then, there exists $M\in\mathbb{N}$ such that for all $N\geq M$, $t\in[0,T]$, the AFAS $\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}$ and $\nu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}$ of Definition 3.40 with initial values $\mu_{0}$ and $\nu_{0}$ satisfy
$\mathcal{W}_{p}(\mu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(t),\nu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(t))\leq\mathcal{W}_{p}(\mu_{0},\nu_{0})e^{\frac{2^{p-3}p\bigl{(}9(p-1)L^{2}+12RL+2R^{2}+3L\bigr{)}+\varepsilon}{p}t}.$ (3.133)
In particular, for every smooth solution $\mu$ of (3.14) with initial value $\mu_{0}$, there exists a smooth solution $\nu$ of (3.14) with initial value $\nu_{0}$ such that
$\mathcal{W}_{p}(\mu(t),\nu(t))\leq\mathcal{W}_{p}(\mu_{0},\nu_{0})e^{\frac{2^{p-3}p\bigl{(}9(p-1)L^{2}+12RL+2R^{2}+3L\bigr{)}}{p}t},$ (3.134)
and vice–versa. Let $S(V,T,\mu_{0})$ be the set of smooth solutions of (3.14) in $\mathcal{W}_{p}(M)$, with initial value $\mu_{0}\in\mathcal{W}_{p}(M)$ on $[0,T]$. In other words, the map
$S:\mathcal{W}_{p}(M)$ $\to\mathbb{P}\biggl{(}\mathcal{C}^{0,\frac{1}{p}}\bigl{(}[0,T],\mathcal{W}_{p}(M)\bigr{)}\biggr{)}$ (3.135)
$\mu$ $\mapsto S(V,T,\mu_{0}),$

### v2

@@ 3.7.1 The linear equation and LAFAS
First, we introduce a linear analogon to (MDEd).
Definition 3.42 (LMDE).
Let $T>0$, $p\geq 2$ and $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $V:[0,T]\to\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$ be continuous with respect to the pp–Wasserstein distance. Let $0,T]\to\mathcal{P}(M)$, $t\mapsto\mu(t)$ be a path of measures. We say that $\mu(t)$ satisfies the linear measure differential equation (LMDE) for $V$ with initial value $\mu_{0}$ if $\mu(0)=\mu_{0}$, $\mu$ is narrowly continuous and for all $s\in[0,T]$ and all $\varphi$ smooth and compactly supported on $M$, we have
$\int_{M}\varphi\operatorname{d}\mu(s)-\int_{M}\varphi\operatorname{d}\mu(0)=\int_{0}^{s}\int_{M}\square^{V_{t}}\varphi\operatorname{d}\mu(t)\operatorname{d}t,$ (LMDE)
@@
$\square^{V_{t}}\varphi=\int_{\mathcal{X}^{2,\infty}(M)}\frac{1}{2}\mathcal{L}_{X-\overline{V_{t}}}^{2}(\varphi)+\mathcal{L}_{X}(\varphi)\operatorname{d}V_{t}(X).$ (3.127)
We introduce an appropriate analogon of the average flow approximation series for (LMDE).
@@
Let $p\geq 2$, $T>0$, $\mu_{0}\in\mathcal{W}_{p}(M)$, $\nu:[0,T]\to\mathcal{W}_{p}(M)$ be continuous with respect to the pp–Wasserstein distance and $V$ be a continuous probability vector–field on $\mathcal{W}_{p}(M)$. Let $\delta>0$ and $\mathcal{}\mathfrak{P}(\delta)$ be a partition of $[0,T]$ with maximal step–length $\mathfrak{p}(\mathcal{}\mathfrak{P}(\delta))\leq\delta$. Let $N\coloneqq|\mathcal{}\mathfrak{P}(\delta)|$ and index the elements of $\mathcal{}\mathfrak{P}(\delta)$ in increasing order. Define
@@
for all $t\in\bigl[x_{l},\frac{x_{l+1}+x_{l}}{2}\bigr]$ and
$\mu^{\mathcal{}\mathfrak{P}(\delta)}(t)=\biggl(e^{V[\nu_{x_{l}}]}_{2\bigl(t-\frac{x_{l+1}+x_{l}}{2}\bigr)}\circ f^{V[\nu_{x_{l}}]}_{x_{l+1}-x_{l}}\biggr)\bigl(\mu^{\mathcal{}\mathfrak{P}(\delta)}(x_{l})\bigr),$
for all $t\in\bigl[\frac{x_{l+1}+x_{l}}{2},x_{l+1}\bigr]$.
@@
Let $p\geq 2$, $T>0$ and $V:[0,T]\to\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$ be continuous. A narrowly continuous curve $\nu:[0,T]\to\mathcal{W}_{p}(M)$ is called a smooth solution of (LMDE) if it solves (LMDE) for $V$ and there exists a sequence $(\mathcal{}\mathfrak{P}_{n})_{n\in\mathbb{N}}$ of partitions of $[0,T]$ with maximal step–length $\mathfrak{p}(\mathcal{}\mathfrak{P}_{n})\to 0$ such that for all $n\in\mathbb{N}$, the LAFAS $\mu^{\mathcal{}\mathfrak{P}_{n}}$ converges pointwise narrowly to $\nu$.
Now, we readily obtain an existence result for equation (LMDE) which is analogous to the Generalised Central Limit Theorem 3.39.
@@
Let $p>2$, $T>0$ and $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $\bigl(\mathcal{}\mathfrak{P}(\frac{1}{N})\bigr)_{N\in\mathbb{N}}$ be a family of partitions of $[0,T]$. Let $\nu:[0,T]\to\mathcal{W}_{p}(M)$ be continuous with respect to the pp–Wasserstein distance and $V$ be a continuous probability vector–field on $\mathcal{W}_{p}(M)$. Assume that $V$ has uniformly bounded pp–th moments, e.g., there exists a constant $B>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $\mathcal{M}_{p}(V[\mu])<B$.
@@
    Assume that $(M,g)$ is of bounded geometry. Assume that there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]\in\mathcal{W}_{p+2}(\mathcal{X}^{2,\infty}(M))$ and $\mathcal{M}_{p+2}(V[\mu])\leq R^{p+2}$. Then, there exists a subsequence $\biggl(\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N_{k}})}\biggr)_{k\in\mathbb{N}}$ of the Linear Average Flow Approximation Series $\biggl(\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\biggr)_{N\in\mathbb{N}}$ with initial value $\mu_{0}$ (See Definition 3.43) which uniformly converges with respect to the pp–Wasserstein distance to a $\frac{1}{p}$–Hölder–continuous solution $\mu:[0,T]\to\mathcal{W}_{p}(M)$ of (LMDE) for $t\mapsto V[\nu(t)]$ with initial value $\mu_{0}$.
@@
    Now, we only assume that there exists $q<p$ such that $V$ is continuous for the $\mathcal{W}_{q}$–distance. Then, there exists a subsequence $\biggl(\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N_{k}})}\biggr)_{k\in\mathbb{N}}$ of the Linear Average Flow Approximation Series with initial value $\mu_{0}$ which converges to a solution $\mu:[0,T]\to\mathcal{W}_{p}(M)$ of (LMDE) for $t\mapsto V[\nu(t)]$ with initial value $\mu_{0}$.
@@ 3.7.2 Grönwall type estimates and Hausdorff–continuity of the set of solutions
Now, we prove two Grönwall type estimates for AFAS and LAFAS respectively for different initial values and identical initial values, but different probability vector–fields.
@@
  • $V$ is a probability vector–field which satisfies:
    • There exists a constant $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]\in\mathcal{W}_{p+2}(\mathcal{X}^{2,\infty}(M))$ and $\mathcal{M}_{p+2}(V[\mu])\leq R^{p+2}$.
@@
Let $\varepsilon>0$. Denote $\mathcal{G}(p,L,R,M)\coloneqq C(M)2^{p-2}\bigl(9(p-1)L^{2}+pR^{2}+6RL+R+3L\bigr)$, where $C(M)$ is a constant depending only on $M$ and $C(\mathbb{R}^{n})=1$ for all $n\in\mathbb{N}$. Then, there exists $M\in\mathbb{N}$ such that for all $N\geq M$, $t\in[0,T]$, the AFAS $\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}$ and $\nu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}$ of Definition 3.43 with initial values $\mu_{0}$ and $\nu_{0}$ satisfy
$\mathcal{W}_{p}(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(t),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(t))\leq\mathcal{W}_{p}(\mu_{0},\nu_{0})e^{\bigl(\mathcal{G}(p,L,R)+\varepsilon\bigr)t}+t\varepsilon.$ (3.128)
Let $p>2$. In particular, for every smooth solution $\mu$ of (MDEd) with initial value $\mu_{0}$, there exists a smooth solution $\nu$ of (MDEd) with initial value $\nu_{0}$ such that
$\mathcal{W}_{p}(\mu(t),\nu(t))\leq\mathcal{W}_{p}(\mu_{0},\nu_{0})e^{\mathcal{G}(p,L,R)t},$ (3.129)
and vice–versa. Let $S_{p}(V,T,\mu_{0})$ be the set of smooth solutions of (MDEd) in $\mathcal{W}_{p}(M)$, with initial value $\mu_{0}\in\mathcal{W}_{p}(M)$ on $[0,T]$. In other words, the map
$S_{p}:\mathcal{W}_{p}(M)$ $\to\mathbb{P}\biggl(\mathcal{C}^{0,\frac{1}{p}}\bigl([0,T],\mathcal{W}_{p}(M)\bigr)\biggr)$ (3.130)
$\mu$ $\mapsto S_{p}(V,T,\mu_{0}),$